40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9258     Entry time: Tue Oct 22 11:58:16 2013
Author: Masayuki 
Type: Update 
Category: Green Locking 
Subject: Yarm ALS PDH 

[Manasa, Masayuki]

Purpose

As this entry, Yarm ALS is not stable enough to lock PRMI + 2 arms. We tried to figure out what is the reason.

What we did

Check connection and alignment

1. Check the Green REFL PD.
Reflection is hitting the center of PD.

2. Check all the BNC connections
All connection are fine.

3. Check which power supply the PDH box is connected to.
PDH box is connected to 1Y4 AC power supply.

Check the control signal and error signal

4. Connected the PZT OUTMON to PC
Before the PZT output was not connected to the monitor channel. We connected that.

5. Saw the time series of the error signal and control signal (PZT output)
 When the Yarm lost end PDH lock, we found that control signal kicked the PZT of end green laser. And also we saw the saturation of control signal. We are not sure where this saturation comes from.

Discussion

With these check, we couldn't find any problem in connection or alignment. But the PDH control signal looks somehow strange. We tried to compare the Yarm signals with that of the Xarm, but we could not conclude anything meaningful.

We don't understand right now but we will continue to check that. We will add more details to the discussion once we have looked into the PDH box signals using oscilloscope.

 

ELOG V3.1.3-